OpenStudy (anonymous):

Last year Leo bought 2 paintings. This year he sold them for $2000 each. On one , he made a 25% profit, and on the other he had a 25% loss. What was his net loss or profit?

OpenStudy (anonymous):

$2000*(0.25) = 500 Profit $2000*(-0.25) = 500 Loss Total profit = 500 - 500 = 0 Total Money = $2000 + $2000 = $4000 Percent Profit = Total profit / total money = 0 / $4000 = 0%

OpenStudy (anonymous):

As he made a 25% profit on one painting, he sold it for 125% of the original price. $2000=125% $1600=100% - its original price was $1600 As he had a 25% loss on the second painting, he sold it for 75% of the original price. $2000=75% $2666.(6)=100% - its original price was $2666.(6) Last year Leo paid for both paintings $1600+$2666.(6)=$4266 and this year he got $4000 for selling them. So, he lost more than $100